# Debiasing Machine Learning Predictions for Causal Inference Without Additional Ground Truth Data: “One Map, Many Trials” in Satellite-Driven Poverty Analysis

### v3

@@ Introduction
Machine learning (ML) models trained on Earth observation (EO) data have recently demonstrated impressive performance in predicting household-level wealth indices, such as the International Wealth Index (IWI), with reported $R^{2}$ values approaching 0.80 (pettersson2023time; jean2016combining; daoud2023using; chi2022microestimates). These EO-ML models offer a much-needed solution to longstanding data scarcity issues in global development, enabling researchers to generate high-resolution, retrospective “wealth maps” in data-poor regions. Indeed, two recurring downstream uses are: (i) evaluating aid or infrastructure programs (e.g., differences in post-intervention wealth across treated vs. control regions), and (ii) tracking regional poverty trends over time.
@@
But while these predictions are promising in the EO-ML upstream task, a key question remains (jerzak2023integrating; jerzak2023image): the value of EO-ML-derived data hinges on whether domain researchers, in the downstream phase, can use these data to estimate causal effects for their specific application areas, from evaluating applied environmental to social science questions (daoudStatisticalModelingThree2023). However, even under ideal conditions, the predictions of EO-ML models tend to have lower variance than the target values, leading to attenuated treatment effects and, thus, potentially misleading conclusions (berglund2012regression). When upstream predictions shrink toward the global mean, estimated effects are biased toward zero, increasing Type-II error: a true 5% effect can appear as only 2–3%, inviting “no impact” conclusions despite meaningful gains. Figure 1 provides a pipeline overview linking the upstream map to multiple downstream trials without requiring new labels.
@@
Recent research has developed methods that can debias upstream machine-learning model predictions, but they require either fresh data in the downstream stage (olofsson2013making; dsl_original; lu2024quantifying) or changing the way the model is trained (ratledge_using_2022). For example, relying on such fresh data, angelopoulos2023prediction propose Prediction-Powered Inference (PPI), a method that corrects any ML model for downstream inference. Similarly, dsl_original achieve the same goal but with better sample efficiency (depieuchon2025benchmarkingdebiasingmethodsllmbased), but it also requires knowing the probability of observing a sample point. While these methods offer a path to integrate upstream and downstream analyses, their success depends on the availability of newly collected treatment-predictor-outcome tuples in the downstream causal step or knowledge of the data labeling process (dsl_original). Often, such data are prohibitively costly or unethical to collect.
@@
ratledge_using_2022 address the challenge of attenuation bias and the unavailability of data in the downstream by adjusting the model’s loss function instead. During training, they penalize quintile-level discrepancies between observed and predicted means, encouraging predictions to match the true outcome distribution more closely across its support. Although their solutions do not need additional downstream data, they require modifying the training pipeline, which is computationally expensive and can degrade predictive performance.
@@
Thus, to enable broader applicability of ML models, the upstream team needs to create a data product that is agnostic to downstream use, requires no downstream-fresh data, produces unbiased causal estimates (due to attenuation), and is computationally efficient (daoudStatisticalModelingThree2023).
@@
We address this gap by proposing a linear calibration method and a family of Tweedie corrections, a novel approach that reduces prediction bias without requiring fresh labeled data. We evaluate these two approaches against three others: Prediction-Powered Inference (PPI), the loss-adjustment approach of ratledge_using_2022, and naive (uncorrected) predictions. The methods introduced here pave the way for the same wealth map to be reused across multiple trials without needing to refit a model for each trial. This also removes the requirement for interaction between the upstream machine-learning team creating the data product and any other domain-knowledge team in the downstream phase. Hence the concept: one map, many trials. We derive analytical proofs and demonstrate the empirical performance of all methods in both simulations and real-world applications, using data from the African Demographic and Health Survey (DHS).
@@
We make these corrections accessible via an open-source software package called unshrink (see Appendix VI); although developed for EO-ML wealth maps, these same attenuation and remedies arise broadly whenever surrogate outcomes replace direct measurements, including in environmental (air pollution, urban heat), demographic (population density), and text/LLM-annotated constructs.
@@
Statistical predictions can systematically shrink estimates toward the central modes of the training distribution (ting2024machine). This phenomenon is rooted in optimal error-minimization estimation (cf. Stein’s paradox), where a model “hedges” predictions in regions of greater uncertainty. When facing noise or limited training data, the optimal strategy, in terms of minimizing residuals, is to pull predictions toward denser regions of the training data. For instance, consider an image that appears with an IWI of 80, but the model has observed this outcome only once, compared to thousands of observations with an IWI of around 70. A perfect prediction would be 80 for this data point, but the model will pull this data point’s prediction towards 70 instead to hedge against the risk of overcommitting to a rare value.
@@
When such biased predictions replace true outcomes in downstream causal analyses, the resulting regression-to-the-mean induces attenuation bias, reducing observed treatment group differences even under randomization (shu2019causal). The difference in mean predicted outcomes between treated and control units shrinks both group means toward the global average, thereby attenuating the estimated effect. In our held-out DHS clusters, the calibration curve exhibits this shrinkage pattern—overpredicting poor locations and underpredicting wealthy ones (Figure 2).
@@
Regression-to-the-mean bias is particularly acute in upstream tasks with scarce ground truth (due to finite samples) or in high-dimensional input data (due to the curse of dimensionality), forcing greater reliance on error-prone predictions—something especially pertinent in development economics, political science, and global health contexts—where household data is often expensive or unavailable—to increasingly substitute upstream predictions (kino2021scoping; kakooei2024mapping; daoudImpactAusterityChildren2024). Moreover, high out-of-sample $R^{2}$ does not eliminate this bias: models can achieve strong accuracy yet yield attenuated causal estimates (ting2024machinelearningmodelssystematically).
@@
Existing remedies are PPI, which subtracts a rectifier estimated from fresh ground-truth pairs at the causal-inference stage (angelopoulos2023prediction), or loss-function adjustments that penalize quintile-level errors (ratledge_using_2022). However, these approaches require that the upstream team adapts the data for downstream analysis, slowing or limiting altogether “one map, many trials” efforts.
@@
To address these limitations, we propose a Tweedie-inspired approach. This Tweedie approach debiases the shrinkage effect in ML-predicted outcomes, enabling downstream researchers to accurately recover causal effects from the upstream data product. Our Tweedie framework treats the ML model as a black box: we assume access only to its out‑of‑sample predictions and do not rely on retraining or modifying the model. Under this setting, we demonstrate that shrinkage can be reversed, either by learning a global calibration correction from held-out labeled data or by estimating a local correction using ideas from Tweedie’s formula, initially described by Robbins and Tweedie in 1956 in the context of empirical Bayes estimation of means under Gaussian observation noise. Both strategies aim to recover unbiased treatment effect estimates from biased predictions, without requiring fresh ground-truth labels at the time of intervention. We benchmark our Tweedie adaptation to existing methods, which are compared in Table A.III.4 based on their associated assumptions and data/training requirements. [^1] As previously mentioned, the only method we are unable to deploy in our setting is the DSL method, as the probability of DHS samples depends on unknown country-level DHS sampling strategies (probabilities not publicly provided at the unit level). In short, Tweedie’s formula is canonical; we adapt it to post-hoc, local de-shrinkage of surrogate outcomes used for causal estimation without new labels, estimating the density score and noise scale from upstream data and applying correction downstream at analysis time.
@@
A key idea here is that the upstream team—a machine-learning team—creates an agnostic dataset (grimmer2022text), meaning that this team’s data product should produce unbiased causal estimates for any number of applications (trials) for a variety of (social science) teams downstream. This setup implies that the upstream team versus the downstream teams should be able to execute on their respective research goals without communication—hence the firewall between the two phases.
@@
In this section, we present a suite of methods designed to debias treatment effect estimates that rely on ML-predicted outcomes. Our overarching goal is to mitigate regression-to-the-mean dynamics in the downstream analysis, improving estimates of causal effects or population-level statistics. We first more formally describe two existing solutions, PPI and the method of ratledge_using_2022, before introducing two no-new-data debiasing approaches—(1) Linear Calibration Correction (LCC) and (2) Tweedie’s Corrections—that aim to correct shrinkage bias without requiring fresh labels in the causal step.
@@
PPI proposes a framework for valid statistical inference when a large set of unlabeled observations is paired with a predictive model’s outputs (angelopoulos2023prediction). Concretely, letting $\theta^{*}$ be the target parameter (e.g., a population mean) and $f$ denote the model that maps covariates $\mathbf{X}_{i}$ to predicted outcomes $f(\mathbf{X}_{i})$, PPI forms an estimator by adding a rectifier term—an adjustment for the average discrepancy between $Y_{i}$ and $f(\mathbf{X}_{i})$ observed in a smaller, labeled dataset. An estimate for $\theta^{*}$ is computed using an “imputed” version based on unlabeled data ($\{\mathbf{X}_{i}\}_{i={n+1}}^{N}$), then corrected by subtracting the bias estimated on the labeled subset. For mean estimates, this adjustment is:
@@
Another approach to remedy attenuation bias is to adjust the loss function: ratledge_using_2022 propose a specialized upstream training procedure that directly penalizes quintile-level discrepancies between observed and predicted means, thereby encouraging $\hat{Y}_{i}$ to match the true distribution of $Y_{i}$ more closely across the support of the outcome distribution. Specifically, the training objective augments the mean squared error loss with an $\ell_{2}$ penalty and a quintile‑bias term:
@@
1. Linear Calibration Correction (LCC)
@@
When $0<k\leq 1$, shrinkage toward the mean is indicated; $m\in\mathbb{R}$ is the intercept, representing the mean fixed bias. The first equality essentially indicates that there is now conditional independence of $\hat{Y}_{i}$ from group/treatment variable $A_{i}\in\{0,1\}$ given $Y_{i}$; the second equality indicates a global affine (linear) calibration of predictions to truth, i.e., $\mathbb{E}[\hat{Y}_{i}\mid Y_{i}]=kY_{i}+m$ with constant slope and intercept. This linearity might not always be empirically satisfied, but appears to occur approximately in our data (see, e.g., Figure A.III.3). When these predictions are used in place of true outcomes to estimate causal effects, this shrinkage attenuates estimated treatment effects.
@@
Proposition 2 (Consistency of the LCC‑corrected estimator)
@@ 2. Tweedie’s correction
However, in ML predictions scenarios, models approximate $\mathbb{E}[Y_{i}\mid\mathbf{X}_{i}]$ for covariates $\mathbf{X}_{i}$ and thus predictions tend to exhibit systematically lower variance than targets due to shrinkage toward the mean (the opposite of what is implied in the baseline classical case, where $\textrm{Var}(\hat{Y}_{i})>\textrm{Var}(Y_{i})$). Indeed, upstream ML models are trained to minimize expected loss over $\mathbf{X}_{i}$ (shalev2014understanding):
@@
where $\theta$ are neural parameters optimized to reduce MSE. Any residual noise $Y_{i}-\hat{Y}_{i}$ thus arises after the prediction, approximately independent of $\hat{Y}_{i}$ itself in distribution (as the model would otherwise exploit correlations to reduce loss). This reasoning therefore inverts the classical view, motivating adoption of the Berkson error model (carroll2006measurement; heid2004two):
@@
To reverse shrinkage under this model without fresh labels, we propose an approach using Tweedie’s formula, introduced by robbins1956empirical and commonly encountered in Gaussian diffusion models:
@@
  • $\mathbb{E}[\omega\mid Z=z]=z+\sigma^{2}\,\frac{d}{dz}\log p_{Z}(z)$   (efron2011tweedie).
@@
Our simulation study benchmarks all debiasing methods and monitors the upstream and downstream analyses while maintaining the firewall between them. The simulation uses a canonical directed acyclic graph (DAG), shown in Figure 3. It is canonical in the sense that the EO-ML literature tends to assume this structure (sakamoto2024scoping), while illustrating the various research phases and corresponding data needs.
In phase 0, the data-generation process (approximating reality) unfolds in the direction of the arrows. In the upstream phase, only the blue-nodes data are available, with finite samples of $Y_{i}$, to produce $\hat{Y_{i}}$ for the entire population $N$. In the downstream, only red nodes are available: $\hat{Y_{i}}$ is now available for $N$, jointly with $C_{i}$ and $A_{is}$. They enable inference on the estimand $\tau_{s}$. (We suppress the index $s$ henceforth.) Appendix IV delineates our simulation further.
@@
We focus on two key objectives in the downstream task: minimizing error and minimizing causal bias. To assess error, we measure how closely the predicted values match the true values of $\tau$, using Mean Absolute Error (MAE). A lower MAE indicates a more accurate recovery of the true treatment effects. To assess bias, we examine the calibration slope from regressing the predicted effects $\hat{\tau}$ on the true $\tau$ values and analyzing the calibration slope. If the estimates are unbiased, the slope should be close to one. Intuitively, a slope of one means predicted and true scales align (no global shrinkage) so downstream treatment effects are not attenuated purely by scaling. Deviations from one indicate systematic over-/underestimation across the range of $\tau$. We report both the slope and its standard error, checking whether the 95% confidence interval includes one. Figure A.III.1 visualizes these slope results.
@@
Here, $Y_{i}$ denotes the International Wealth Index (IWI), a continuous asset-based wealth score (0–100) aggregated at the cluster level to mitigate noise (burkeUsingSatelliteImagery2021; pmlr-v275-zhu25a). The covariate array $\mathbf{X}_{i}$ consists of Landsat satellite imagery (30 m resolution, six multispectral bands) centered on each DHS cluster (6.72 km × 6.72 km tiles), processed by taking the three-year median over cloud-free pixels leading up to the survey date.
Two models were fitted for each fold, one using MSE and one using the loss introduced by ratledge_using_2022. The models shared the same ResNet-50 architecture and were initialized with weights pretrained on the SSL4EO-L dataset (stewart2023ssl4eoldatasetsfoundationmodels). For more details about model training, see Appendix II. The resulting MSE models all have an $R^{2}$ score between 0.74 and 0.76 on their hold-out fold, and an average slope of 0.79 . Their combined predictions can be seen in Figure 2. The Ratledge models have a slope of 0.99, but at the cost of reduced $R^{2}$ scores between 0.67 and 0.70.
@@
As expected, the MSE of the naive model exhibits shrinkage behavior: it tends to overpredict the wealth of poorer regions and underpredict the wealth of richer regions. As shown in Figure 4, Tweedie’s correction accurately compensates for this, not only resulting in less biased estimates (regression slope of 0.83 to 0.90), but also decreases the MAE from 2.67 to 2.39. In fact, if we consider only regions with at least 100 clusters, the Tweedie-corrected estimates exhibit virtually no shrinkage bias, with a slope of 0.99. Appendix III explicates.
@@
To assess the practical utility of our correction methods in a setting where the true treatment effect is unknown, we construct an evaluation framework based on real-world development projects. Specifically, we leverage the same dataset of geo-referenced interventions funded by the World Bank and the Chinese government as was used in (malik2021banking; Conlin2024). Each project is tagged with an associated aid sector $s\in\mathcal{S}$ (e.g., Health, Water Supply and Sanitation, or Women in Development), which defines a downstream trial $D_{\text{Trial }s}=\{(\mathbf{X}_{i},\hat{Y}_{i},A_{is}):i\in\mathcal{I}_{\textrm{Trial }s}\}$. Here, $A_{is}\in\{0,1\}$ indicates exposure to a sector-$s$ intervention at location $i$, $\mathbf{X}_{i}$ are satellite features, and $\hat{Y}_{i}$ are IWI predictions from the upstream-trained model. Figure A.III.2 depicts the spatial distribution of treatments across the continent.
@@
See Figure 5 for results. Each point is a causal study. While the $x$-axis shows what would have happened when using the label $Y_{i}$ in these studies, the $y$-axis shows the corresponding causal effect using $\hat{Y_{i}}$ with its corresponding adjustment approach. Orange lines check the correlation among these data. A score of one is perfect. The analysis shows that Tweedie’s correction (panel d) comes close to this desiderata, with a correlation of $0.998$; and it does so with the lowest MAE of $0.57$. Had we used the naive predictions (panel a), the correlation would be $0.958$ and the MAE would be $0.69$. Assuming the PPI set up, with 10% fresh data available, we would have achieved a $0.981$; and it does so with the lowest MAE of $0.73$. The PPI performance is, however, a function of available data in the downstream, and panel f shows that with increasing fresh data, the PPI correction will outperform our Tweedie. However, since PPI demands substantial amounts of such data, this method would necessitate costly new surveys, resulting in large expenditures for downstream teams (depieuchon2025benchmarkingdebiasingmethodsllmbased).
@@
That said, bias can remain or be reintroduced when (a) the joint distribution of $(\mathbf{X},Y)$ shifts between up and downstream, (b) errors are strongly heteroskedastic or non-Gaussian so that a single $\sigma$ over- or under-corrects in parts of the support, or (c) $\hat{\sigma}$ is misestimated (e.g., from overfitting late in training). We therefore (i) recommend estimating $\sigma$ on the training split to reduce over-correction, (ii) report block-bootstrap uncertainty for spatial analyses, and (iii) monitor calibration after correction.
For future research, we see three immediate paths. First, formalize a theory of adaptive corrections that locally tunes the strength of Tweedie’s adjustment. Second, extend the methodology to multi-task settings where outcomes are predicted jointly. Third, integrate causal discovery into the pipeline to co-model confounding and attenuation. $\square$
Table of Contents for Appendices
Appendix I. Deriving Implications of Tweedie’s Formula for Causal Debiasing
Appendix II. Neural Model Details
Appendix III. Additional Empirical Results
Appendix IV. Simulation Design
Appendix V. World Bank and China Empirical Case
Appendix VI. Implementing Debiasing Strategies with unshrink
Appendix I. Deriving Implications of Tweedie’s Formula for Causal Debiasing

### v4

@@ Introduction
Machine learning (ML) models trained on Earth observation (EO) data have recently demonstrated impressive performance in predicting household-level wealth indices, such as the International Wealth Index (IWI), with reported $R^{2}$ values approaching 0.80 (Pettersson et al. 2023; Jean et al. 2016; Daoud et al. 2023; Chi et al. 2022). These EO-ML models offer a much-needed solution to longstanding data scarcity issues in global development, enabling researchers to generate high-resolution, retrospective “wealth maps” in data-poor regions. Indeed, two recurring downstream uses are: (i) evaluating aid or infrastructure programs (e.g., differences in post-intervention wealth across treated vs. control regions), and (ii) tracking regional poverty trends over time.
@@
But while these predictions are promising in the EO-ML upstream task, a key question remains (Jerzak, Johansson, and Daoud 2023a; Jerzak, Johansson, and Daoud 2023b): the value of EO-ML-derived data hinges on whether domain researchers, in the downstream phase, can use these data to estimate causal effects for their specific application areas, from evaluating applied environmental to social science questions (Daoud and Dubhashi 2023). However, even under ideal conditions, the predictions of EO-ML models tend to have lower variance than the target values, leading to attenuated treatment effects and, thus, potentially misleading conclusions (Berglund 2012). When upstream predictions shrink toward the global mean, estimated effects are biased toward zero, increasing Type-II error: a true 5% effect can appear as only 2–3%, inviting “no impact” conclusions despite meaningful gains. Figure 1 provides a pipeline overview linking the upstream map to multiple downstream trials without requiring new labels.
@@
Recent research has developed methods that can debias upstream machine-learning model predictions, but they require either fresh data in the downstream stage (Olofsson et al. 2013; Egami et al. 2023; Lu, Bates, and Wang 2024) or changing the way the model is trained (Ratledge et al. 2022). For example, relying on such fresh data, Angelopoulos et al. (2023) propose Prediction-Powered Inference (PPI), a method that corrects any ML model for downstream inference. Similarly, Egami et al. (2023) achieve the same goal but with better sample efficiency (de Pieuchon et al. 2025), but it also requires knowing the probability of observing a sample point. While these methods offer a path to integrate upstream and downstream analyses, their success depends on the availability of newly collected treatment-predictor-outcome tuples in the downstream causal step or knowledge of the data labeling process (Egami et al. 2023). Often, such data are prohibitively costly or unethical to collect.
@@
Ratledge et al. (2022) address the challenge of attenuation bias and the unavailability of data in the downstream by adjusting the model’s loss function instead. During training, they penalize quintile-level discrepancies between observed and predicted means, encouraging predictions to match the true outcome distribution more closely across its support. Although their solutions do not need additional downstream data, they require modifying the training pipeline, which is computationally expensive and can degrade predictive performance.
@@
Thus, to enable broader applicability of ML models, the upstream team needs to create a data product that is agnostic to downstream use, requires no downstream-fresh data, produces unbiased causal estimates (due to attenuation), and is computationally efficient (Daoud and Dubhashi 2023).
@@
We address this gap by proposing a linear calibration method and a family of Tweedie corrections, a novel approach that reduces prediction bias without requiring fresh labeled data. We evaluate these two approaches against three others: Prediction-Powered Inference (PPI), the loss-adjustment approach of Ratledge et al. (2022), and naive (uncorrected) predictions. The methods introduced here pave the way for the same wealth map to be reused across multiple trials without needing to refit a model for each trial. This also removes the requirement for interaction between the upstream machine-learning team creating the data product and any other domain-knowledge team in the downstream phase. Hence the concept: one map, many trials. We derive analytical proofs and demonstrate the empirical performance of all methods in both simulations and real-world applications, using data from the African Demographic and Health Survey (DHS).
@@
We make these corrections accessible via an open-source software package called unshrink (see Appendix F); although developed for EO-ML wealth maps, these same attenuation and remedies arise broadly whenever surrogate outcomes replace direct measurements, including in environmental (air pollution, urban heat), demographic (population density), and text/LLM-annotated constructs.
@@
Statistical predictions can systematically shrink estimates toward the central modes of the training distribution (Ting 2024a). This phenomenon is rooted in optimal error-minimization estimation (cf. Stein’s paradox), where a model “hedges” predictions in regions of greater uncertainty. When facing noise or limited training data, the optimal strategy, in terms of minimizing residuals, is to pull predictions toward denser regions of the training data. For instance, consider an image that appears with an IWI of 80, but the model has observed this outcome only once, compared to thousands of observations with an IWI of around 70. A perfect prediction would be 80 for this data point, but the model will pull this data point’s prediction towards 70 instead to hedge against the risk of overcommitting to a rare value.
@@
When such biased predictions replace true outcomes in downstream causal analyses, the resulting regression-to-the-mean induces attenuation bias, reducing observed treatment group differences even under randomization (Shu and Yi 2019). The difference in mean predicted outcomes between treated and control units shrinks both group means toward the global average, thereby attenuating the estimated effect. In our held-out DHS clusters, the calibration curve exhibits this shrinkage pattern—overpredicting poor locations and underpredicting wealthy ones (Figure 2).
@@
Regression-to-the-mean bias is particularly acute in upstream tasks with scarce ground truth (due to finite samples) or in high-dimensional input data (due to the curse of dimensionality), forcing greater reliance on error-prone predictions—something especially pertinent in development economics, political science, and global health contexts—where household data is often expensive or unavailable—to increasingly substitute upstream predictions (Kino et al. 2021; Kakooei et al. 2024; Daoud and Johansson 2024). Moreover, high out-of-sample $R^{2}$ does not eliminate this bias: models can achieve strong accuracy yet yield attenuated causal estimates (Ting 2024b).
@@
Existing remedies are PPI, which subtracts a rectifier estimated from fresh ground-truth pairs at the causal-inference stage (Angelopoulos et al. 2023), or loss-function adjustments that penalize quintile-level errors (Ratledge et al. 2022). However, these approaches require that the upstream team adapts the data for downstream analysis, slowing or limiting altogether “one map, many trials” efforts.
@@
To address these limitations, we propose a Tweedie-inspired approach. This Tweedie approach debiases the shrinkage effect in ML-predicted outcomes, enabling downstream researchers to accurately recover causal effects from the upstream data product. Our Tweedie framework treats the ML model as a black box: we assume access only to its out‑of‑sample predictions and do not rely on retraining or modifying the model. Under this setting, we demonstrate that shrinkage can be reversed, either by learning a global calibration correction from held-out labeled data or by estimating a local correction using ideas from Tweedie’s formula, initially described by Robbins and Tweedie in 1956 in the context of empirical Bayes estimation of means under Gaussian observation noise. Both strategies aim to recover unbiased treatment effect estimates from biased predictions, without requiring fresh ground-truth labels at the time of intervention. We benchmark our Tweedie adaptation to existing methods, which are compared in Table 5 based on their associated assumptions and data/training requirements. [^1] As previously mentioned, the only method we are unable to deploy in our setting is the DSL method, as the probability of DHS samples depends on unknown country-level DHS sampling strategies (probabilities not publicly provided at the unit level). In short, Tweedie’s formula is canonical; we adapt it to post-hoc, local de-shrinkage of surrogate outcomes used for causal estimation without new labels, estimating the density score and noise scale from upstream data and applying correction downstream at analysis time.
@@
A key idea here is that the upstream team—a machine-learning team—creates an agnostic dataset (Grimmer, Roberts, and Stewart 2022), meaning that this team’s data product should produce unbiased causal estimates for any number of applications (trials) for a variety of (social science) teams downstream. This setup implies that the upstream team versus the downstream teams should be able to execute on their respective research goals without communication—hence the firewall between the two phases.
@@
In this section, we present a suite of methods designed to debias treatment effect estimates that rely on ML-predicted outcomes. Our overarching goal is to mitigate regression-to-the-mean dynamics in the downstream analysis, improving estimates of causal effects or population-level statistics. We first more formally describe two existing solutions, PPI and the method of Ratledge et al. (2022), before introducing two no-new-data debiasing approaches—(1) Linear Calibration Correction (LCC) and (2) Tweedie’s Corrections—that aim to correct shrinkage bias without requiring fresh labels in the causal step.
@@
PPI proposes a framework for valid statistical inference when a large set of unlabeled observations is paired with a predictive model’s outputs (Angelopoulos et al. 2023). Concretely, letting $\theta^{*}$ be the target parameter (e.g., a population mean) and $f$ denote the model that maps covariates $\mathbf{X}_{i}$ to predicted outcomes $f(\mathbf{X}_{i})$, PPI forms an estimator by adding a rectifier term—an adjustment for the average discrepancy between $Y_{i}$ and $f(\mathbf{X}_{i})$ observed in a smaller, labeled dataset. An estimate for $\theta^{*}$ is computed using an “imputed” version based on unlabeled data ($\{\mathbf{X}_{i}\}_{i={n+1}}^{N}$), then corrected by subtracting the bias estimated on the labeled subset. For mean estimates, this adjustment is:
@@
Another approach to remedy attenuation bias is to adjust the loss function: Ratledge et al. (2022) propose a specialized upstream training procedure that directly penalizes quintile-level discrepancies between observed and predicted means, thereby encouraging $\hat{Y}_{i}$ to match the true distribution of $Y_{i}$ more closely across the support of the outcome distribution. Specifically, the training objective augments the mean squared error loss with an $\ell_{2}$ penalty and a quintile‑bias term:
@@
Linear Calibration Correction (LCC)
@@
When $0<k\leq 1$, shrinkage toward the mean is indicated; $m\in\mathbb{R}$ is the intercept, representing the mean fixed bias. The first equality essentially indicates that there is now conditional independence of $\hat{Y}_{i}$ from group/treatment variable $A_{i}\in\{0,1\}$ given $Y_{i}$; the second equality indicates a global affine (linear) calibration of predictions to truth, i.e., $\mathbb{E}[\hat{Y}_{i}\mid Y_{i}]=kY_{i}+m$ with constant slope and intercept. This linearity might not always be empirically satisfied, but appears to occur approximately in our data (see, e.g., Figure 9). When these predictions are used in place of true outcomes to estimate causal effects, this shrinkage attenuates estimated treatment effects.
@@
Proposition 2 (Consistency of LCC correction)
@@
However, in ML predictions scenarios, models approximate $\mathbb{E}[Y_{i}\mid\mathbf{X}_{i}]$ for covariates $\mathbf{X}_{i}$ and thus predictions tend to exhibit systematically lower variance than targets due to shrinkage toward the mean (the opposite of what is implied in the baseline classical case, where $\textrm{Var}(\hat{Y}_{i})>\textrm{Var}(Y_{i})$). Indeed, upstream ML models are trained to minimize expected loss over $\mathbf{X}_{i}$ (Shalev-Shwartz and Ben-David 2014):
@@
where $\theta$ are neural parameters optimized to reduce MSE. Any residual noise $Y_{i}-\hat{Y}_{i}$ thus arises after the prediction, approximately independent of $\hat{Y}_{i}$ itself in distribution (as the model would otherwise exploit correlations to reduce loss). This reasoning therefore inverts the classical view, motivating adoption of the Berkson error model (Carroll et al. 2006; Heid et al. 2004):
@@
To reverse shrinkage under this model without fresh labels, we propose an approach using Tweedie’s formula, introduced by Robbins (1956) and commonly encountered in Gaussian diffusion models:
@@
  • $\mathbb{E}[\omega\mid Z=z]=z+\sigma^{2}\,\frac{d}{dz}\log p_{Z}(z)$   (Efron 2011).
@@
Our simulation study benchmarks all debiasing methods and monitors the upstream and downstream analyses while maintaining the firewall between them. The simulation uses a canonical directed acyclic graph (DAG), shown in Figure 3. It is canonical in the sense that the EO-ML literature tends to assume this structure (Sakamoto, Jerzak, and Daoud 2025), while illustrating the various research phases and corresponding data needs.
In phase 0, the data-generation process (approximating reality) unfolds in the direction of the arrows. In the upstream phase, only the blue-nodes data are available, with finite samples of $Y_{i}$, to produce $\hat{Y_{i}}$ for the entire population $N$. In the downstream, only red nodes are available: $\hat{Y_{i}}$ is now available for $N$, jointly with $C_{i}$ and $A_{is}$. They enable inference on the estimand $\tau_{s}$. (We suppress the index $s$ henceforth.) Appendix D delineates our simulation further.
@@
We focus on two key objectives in the downstream task: minimizing error and minimizing causal bias. To assess error, we measure how closely the predicted values match the true values of $\tau$, using Mean Absolute Error (MAE). A lower MAE indicates a more accurate recovery of the true treatment effects. To assess bias, we examine the calibration slope from regressing the predicted effects $\hat{\tau}$ on the true $\tau$ values and analyzing the calibration slope. If the estimates are unbiased, the slope should be close to one. Intuitively, a slope of one means predicted and true scales align (no global shrinkage) so downstream treatment effects are not attenuated purely by scaling. Deviations from one indicate systematic over-/underestimation across the range of $\tau$. We report both the slope and its standard error, checking whether the 95% confidence interval includes one. Figure 7 visualizes these slope results.
@@
Here, $Y_{i}$ denotes the International Wealth Index (IWI), a continuous asset-based wealth score (0–100) aggregated at the cluster level to mitigate noise (Burke et al. 2021; Zhu, Jerzak, and Daoud 2025). The covariate array $\mathbf{X}_{i}$ consists of Landsat satellite imagery (30 m resolution, six multispectral bands) centered on each DHS cluster (6.72 km × 6.72 km tiles), processed by taking the three-year median over cloud-free pixels leading up to the survey date.
Two models were fitted for each fold, one using MSE and one using the loss introduced by Ratledge et al. (2022). The models shared the same ResNet-50 architecture and were initialized with weights pretrained on the SSL4EO-L dataset (Stewart et al. 2023). For more details about model training, see Appendix B. The resulting MSE models all have an $R^{2}$ score between 0.74 and 0.76 on their hold-out fold, and an average slope of 0.79 . Their combined predictions can be seen in Figure 2. The Ratledge models have a slope of 0.99, but at the cost of reduced $R^{2}$ scores between 0.67 and 0.70.
@@
As expected, the MSE of the naive model exhibits shrinkage behavior: it tends to overpredict the wealth of poorer regions and underpredict the wealth of richer regions. As shown in Figure 4, Tweedie’s correction accurately compensates for this, not only resulting in less biased estimates (regression slope of 0.83 to 0.90), but also decreases the MAE from 2.67 to 2.39. In fact, if we consider only regions with at least 100 clusters, the Tweedie-corrected estimates exhibit virtually no shrinkage bias, with a slope of 0.99. Appendix C explicates.
@@
To assess the practical utility of our correction methods in a setting where the true treatment effect is unknown, we construct an evaluation framework based on real-world development projects. Specifically, we leverage the same dataset of geo-referenced interventions funded by the World Bank and the Chinese government as was used in (Malik et al. 2021; Conlin 2024). Each project is tagged with an associated aid sector $s\in\mathcal{S}$ (e.g., Health, Water Supply and Sanitation, or Women in Development), which defines a downstream trial $D_{\text{Trial }s}=\{(\mathbf{X}_{i},\hat{Y}_{i},A_{is}):i\in\mathcal{I}_{\textrm{Trial }s}\}$. Here, $A_{is}\in\{0,1\}$ indicates exposure to a sector-$s$ intervention at location $i$, $\mathbf{X}_{i}$ are satellite features, and $\hat{Y}_{i}$ are IWI predictions from the upstream-trained model. Figure 8 depicts the spatial distribution of treatments across the continent.
@@
See Figure 5 for results. Each point is a causal study. While the $x$-axis shows what would have happened when using the label $Y_{i}$ in these studies, the $y$-axis shows the corresponding causal effect using $\hat{Y_{i}}$ with its corresponding adjustment approach. Orange lines check the correlation among these data. A score of one is perfect. The analysis shows that Tweedie’s correction (panel d) comes close to this desiderata, with a correlation of $0.998$; and it does so with the lowest MAE of $0.57$. Had we used the naive predictions (panel a), the correlation would be $0.958$ and the MAE would be $0.69$. Assuming the PPI set up, with 10% fresh data available, we would have achieved a $0.981$; and it does so with the lowest MAE of $0.73$. The PPI performance is, however, a function of available data in the downstream, and panel f shows that with increasing fresh data, the PPI correction will outperform our Tweedie. However, since PPI demands substantial amounts of such data, this method would necessitate costly new surveys, resulting in large expenditures for downstream teams (de Pieuchon et al. 2025).
@@
That said, bias can remain or be reintroduced when (a) the joint distribution of $(\mathbf{X},Y)$ shifts between up and downstream, (b) errors are strongly heteroskedastic or non-Gaussian so that a single $\sigma$ over- or under-corrects in parts of the support, or (c) $\hat{\sigma}$ is misestimated (e.g., from overfitting late in training). We therefore (i) recommend estimating $\sigma$ on the training split to reduce over-correction, (ii) report block-bootstrap uncertainty for spatial analyses, and (iii) monitor calibration after correction. $\square$
Appendix A A. Deriving Implications of Tweedie’s Formula for Causal Debiasing